A generic salt, AB3, has a molar mass of 185 g/mol and a solubility of 5.10 g/L at 25 °C. What is the Ksp of this salt at 25 °C?
Answer #1

AB3---->A+3 +3B-

Ksp = [A+3] [B-]3

Solubility = 5.1 g/L/ 185 g/mol = 0.028 gmoles/L

forom the equation [A+3] = 0.028 and [B-] =3*0.028= 0.084

KS[ = 0.028*(0.084)3=1.66*10-5